[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Sep 06, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Sep 7 02:29:40 UTC 2016


HSU Shift Log for Tue Sep 06 2016 22:29:32 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: None

Other Techincal issues: Black toner is very low in the HSU ops printer. A  
NCO helpdesk ticket was submitted to have it replaced.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160907/068bb2a1/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list